### v1

# "I understand why I got this grade": Automatic Short Answer Grading with Feedback

Abstract
The demand for efficient and accurate assessment methods has intensified as education systems transition to digital platforms. Providing feedback is essential in educational settings and goes beyond simply conveying marks as it justifies the assigned marks. In this context, we present a significant advancement in automated grading by introducing Engineering Short Answer Feedback (EngSAF)—a dataset of student answers accompanied by reference answers and questions for the Automatic Short Answer Grading (ASAG) task. The EngSAF dataset is meticulously curated to cover a diverse range of subjects, questions, and answer patterns from multiple engineering domains. We leverage state-of-the-art large language models’ (LLMs) generative capabilities with our Label-Aware Synthetic Feedback Generation (LASFG) strategy to include feedback in our dataset. This paper underscores the importance of enhanced feedback in practical educational settings, outlines dataset annotation and feedback generation processes, conducts a thorough EngSAF analysis, and provides different LLMs-based zero-shot and finetuned baselines for future comparison111Our code and dataset are available at https://github.com/dishankaggarwal/EngSAF. Additionally, we demonstrate the efficiency and effectiveness of the ASAG system through its deployment in a real-world end-semester exam at the Indian Institute of Technology Bombay (IITB), showcasing its practical viability and potential for broader implementation in educational institutions.
1 Introduction
Technology integration in education has resulted in transformative changes, redefining traditional pedagogical approaches and assessment methodologies. Effective education relies on feedback and explanations provided during assessments to ensure quality learning outcomes Shute (2008). Grading questions in tests and examinations have proven to be a good measure to assess student’s learning and understanding of a topic or a subject. An exam could include various question types, such as multiple choice, fill-in-the-blank, short answers, essays, etc. Among these question types, short answers and essays are more complicated to analyze than multiple-choice or fill-in-the-blank type questions due to flexibility and natural language in the response. Automating the grading process becomes crucial, especially in countries with extremely high student-to-teacher ratios, as it can significantly reduce instructor’s workloads and improve the assessment process. Significant advancements have been achieved in this field in recent years, primarily after the introduction of SemEval Semantic Textual Similarity (STS) task Agirre et al. (2012, 2013, 2014, 2015).
This challenge can be approached as a machine learning issue, where the objective is to grade a student’s response based on how similar it is to the reference answers. However, simply assigning a score or label to a learner’s response is often inadequate in practical educational contexts. Ahea et al. (2016) highlights the value and effectiveness of feedback in improving students’ learning and professionalizing teaching in higher education.
1.1 Problem Statement
Given a question, a reference answer, and a student’s answer, the aim is to provide content-focused elaborated feedback and assign a label indicating the degree of correctness in the student’s answer compared to the reference answer. Here, we focus on questions where the answer type is a sentence or a short paragraph. This task involves evaluating the alignment between the student’s and reference answers. Figure 1 illustrates the problem statement, showing the input and output using an example from the EngSAF dataset.
| Question | What is the difference between basin order and channel order? |
| Reference Answer | Basin order is highest order of any stream in that basin whereas channel order is order of stream which denotes that in what order of streams has joined the channel." |
| Student Answer 1 | Highest order channel is the basin order whereas channel order is the order of channel from tributaries to reaches to main river stream. |
| Label | 2 (Correct response) |
| Feedback | Excellent! You have a clear understanding of the distinction between basin order and channel order. |
| Student Answer 2 | Channel order reflects to the number of streams coming together to form a channel. |
| Label | 1 (Partially Correct response) |
| Feedback | Your answer includes a part of the distinction. Channel order indeed indicates the number of streams joining together to make a channel, but the difference between basin order and channel order is not mentioned. |
| Student Answer 3 | Channel order is the order of the highest order streams. For example, two first order streams (or more) will make a second order stream and similarly for highest orders. |
| Label | 0 (Incorrect response) |
| Feedback | The student answer confuses basin order with channel order. Basin order refers to the highest order of streams within a basin, while channel order refers to the order of streams based on the sequence of junctions. |
1.2 Motivation
The increasing demand for technology in education has led to a need for more efficient and effective methods of grading short-answer assessments. In the case of short answer grading, feedback with clear explanations goes beyond simply conveying a grade. It offers valuable insights into student’s strengths, weaknesses, and areas of improvement. However, the effectiveness of feedback is largely unexplored due to the lack of public, content-centered, elaborated feedback datasets in different domains. These datasets are crucial for training and developing automated feedback systems that can provide personalized and nuanced feedback. The need for a more efficient and effective ASAG system incorporating feedback alongside grade has arisen due to these limitations, and this is where the Engineering Short Answer Feedback (EngSAF) dataset steps in, consisting of questions students answer from multiple engineering domains for the ASAG task.
Our Contributions are:
1.
EngSAF dataset containing around 5.8K student responses to 119 questions from multiple engineering domains along with synthetically generated feedback explaining the assigned grade for the task of ASAG. To the best of our knowledge, this is the first dataset containing questions and responses from multiple engineering domains. (Section 3)
2.
Benchmark scores on the EngSAF dataset using different Large Language Models (LLM) for future comparison and research. (Section 5)
3.
Real-world deployment of the EngSAF fine-tuned ASAG model in an end-semester exam at Indian Institute of Technology, Bombay (IITB) (Section 6).
2 Related work
2.1 Automatic Short Answer Grading (ASAG)
ASAG is an essential area of research that has garnered significant attention in recent years. Several approaches have been proposed for traditional ASAG, ranging from rule-based methods to more sophisticated machine-learning techniques. One early approach for ASAG was based on keyword or pattern matching, where the presence or absence of certain keywords in the student’s answer was used to determine its accuracy. Mitchell et al. (2002); Sukkarieh et al. (2004); Nielsen et al. (2009).To overcome these limitations, researchers have developed more sophisticated methods that use natural language processing (NLP) techniques. One such method is based on Latent Semantic Analysis (LSA), which represents texts as high-dimensional vectors and compares them to the reference answers using cosine similarity LaVoie et al. (2020). In a related study, the task of ASAG is addressed by incorporating features such as answer length, grammatical correctness, and semantic similarity in comparison to reference answers Sultan et al. (2016).
More recently, deep learning models such as convolutional neural networks (CNNs) and recurrent neural networks (RNNs) have been applied to ASAG’s task. These models are trained on large amounts of annotated data and can capture the semantic relationships between words in a student’s answer, and the reference answers Surya et al. (2019); Zhang et al. (2022). Pre-trained Language Models (PLMs), such as BERT Devlin et al. (2018), GPT Radford et al. (2019), RoBERTa Liu et al. (2019), DistillBERT Sanh et al. (2019), and ALBERT Lan et al. (2019) have performed exceptionally well for various tasks in NLP, also used for the task of ASAG. Sentence-BERT or SBERT Reimers and Gurevych (2019) performs exceptionally well on current publicly available traditional ASAG datasets Condor et al. (2021). A common limitation of the prior work is that only grades are assigned, and no feedback is provided. We seek to fill this gap with the introduction of the EngSAF dataset.
2.2 Short Answer Grading Datasets
Several publicly accessible datasets have been curated, each designed to facilitate research and benchmarking in the ASAG task. In 2009, Mohler and Milhalcea published a short answer dataset on a data structures course that contained 630 records Mohler and Mihalcea (2009). In 2011, they published an extended dataset on the same course that contained 2273 records Mohler et al. (2011), named the Texas Extended dataset. To complete the assignment of student-authored answer annotation, the Beetle corpus is being created Nielsen et al. (2008). Another dataset that is publicly available on Kaggle and could be used for the task of ASAG is provided by Hewlett Foundation named ASAP-AES222https://www.kaggle.com/c/asap-aes/data (Automated Assessment Prize Competition for Essay Scoring).
However, all aforementioned datasets are conventional ASAG datasets solely consisting of grades or scores. Our extensive literature survey includes only one short answer grading dataset with content-focused elaborated feedback, i.e., Filighera et al. (2022), which introduces an inaugural dataset for short-answer feedback comprising bilingual responses in English and German. Human annotations were meticulously collected and refined to uphold the feedback’s quality. However, this dataset is constrained by its limited number of student responses () in English and its exclusive focus on questions from a single domain, specifically computer network queries, thereby lacking diversity across different engineering domains. To address these limitations, we introduced EngSAF, which contains data points and questions from multiple engineering domains.
3 Engineering Short Answer Feedback (EngSAF) Dataset
This dataset contains 119 questions drawn from different undergraduate and graduate engineering courses, accompanied by approximately 5.8k student responses. An Instructor provided correct answer/ reference answer also accompanies each question. These questions and answers have been taken from actual quizzes/ exams from the Indian Institute of Technology, Bombay (IITB), covering a diverse range of 25 courses. The questions and responses span across multiple domains, including image processing, water quality management, and operating systems, to name a few. Technical Appendix contains the complete list of courses included in the dataset. Based on the instructor’s assigned marks, each response has been categorized as “correct,” “partially correct,” or “incorrect.”
Table 1 provides an example from the EngSAF dataset showcasing a question, its corresponding reference answer, and three different student responses alongside their associated output labels and synthetically generated feedback/Explanation. Technical Appendix contains more examples from the EngSAF dataset. The dataset is applicable for both traditional automatic short answer grading and the generation of elaborated feedback.
3.1 Dataset Construction
EngSAF contains questions and student responses from different undergraduate and graduate engineering courses. The instruction provided a correct answer/ reference answer accompanying each question. The Instructor/ Teaching Assistant (TA) assigns marks to each student’s answer from their respective course. The students’ responses and grades are obtained from a reputed engineering university. The quality assessment of the proposed dataset is covered in subsection 3.5. If the maximum mark for a question is and a student’s answer is graded marks by the instructor, then the output label is annotated according to the scheme in Table 2.
| Condition | Output Label |
| 2 (Correct Response) | |
| 1 (Partially Correct Response) | |
| 0 (Incorrect Response) |
3.2 Challenges and Requirements
Diversity in student responses: Human language is inherently subjective, and different individuals may interpret the same content differently. This subjectivity extends to the grading process, making it challenging to provide universally applicable feedback.
Response Variability: Short answer responses can vary significantly regarding quality, coherence, and relevance. Some responses may be well-structured and articulate, while others may be incomplete or contain grammatical errors. Designing feedback that caters to this variability in response quality poses a considerable challenge.
Feedback Impact: Wrong feedback can negatively shape learners’ perceptions of themselves and their abilities. If the feedback consistently highlights their shortcomings or mistakes, they may internalize these negative perceptions and develop a fixed mindset about their capabilities. This can hinder their willingness to take risks, seek challenges, and persist in facing difficulties.
3.3 Label Aware Synthetic Feedback Generation (LASFG)
Leveraging the advanced language generative and reasoning capabilities of state-of-the-art Large Language Models (LLMs) like Gemini333https://gemini.google.com/ and ChatGPT, we enhance educational assessment with feedback, particularly in Automatic Short Answer Grading (ASAG). The approach involves utilizing Gemini’s ability to comprehend input prompts consisting of a question, a student’s answer, a reference answer, and the corresponding grading label provided by an instructor or teaching assistant to generate content-focused elaborated feedback. The generated feedback covers the reasoning or explanation of the Gold output label. Grammarly444https://www.grammarly.com/ then checks the synthetically generated feedback to remove any grammatical errors. Prompt details used for synthetic feedback generation can be found in the Technical Appendix 10.1. The quality estimation of the generated feedback is shown in Section 3.5
3.4 Corpus Statistics
| Label | Train | UA | UQ | Total |
| Correct | 1716 | 403 | 321 | 2440 |
| Partially Correct | 1412 | 344 | 278 | 2034 |
| Incorrect | 941 | 233 | 166 | 1340 |
| Total | 4069 | 980 | 765 | 5814 |
Following Dzikovska et al. (2013), we partitioned the data into training sets, comprising 70% of the dataset, as well as unseen answers (16%) and unseen questions (14%) for test sets as shown in table 3. The test split with unseen answers includes fresh responses to the questions used in training. In contrast, the test split with unseen questions comprises entirely new questions for testing the model’s ability to generalize to new questions without prior exposure. Technical Appendix 10.3 contains insight about insights about the distribution of text length for student answers, reference answers, questions, and feedback in the EngSAF dataset.
3.5 Quality Estimation
To show the reliability and credibility of our dataset and synthetically generated feedback, we randomly sampled 300 data points and equally distributed them across the output label. This sampled data is a good representation of the EngSAF dataset.
Each synthetically generated feedback was scored by three human annotators over three aspects. Each aspect is scored on a scale (1-5), with a high score indicating a better response. Each feedback is analyzed on the following aspects.
1.
Fluency and Grammatical Correctness: This aspect tests whether the generated feedback is fluent in English and grammatically correct.
2.
Feedback Correctness/Accuracy: This aspect assesses the overall quality of the generated feedback regarding content, relevance, quality, and explanation for the assigned grade.
3.
Emotional Impact: This aspect assesses how feedback affects the learner’s emotional state. Annotators are tasked with rating whether the feedback avoids triggering negative emotions or impacts by refraining from using words such as "fail" that may evoke feelings of discouragement or distress in the learner.
| Aspect | Avg. Score |
| Fluency & Grammatical Correctness | 4.73 |
| Feedback Correctness/Accuracy | 4.55 |
| Emotional Impact | 4.61 |
Three human annotators evaluate the correctness of each output label for sampled data points, achieving an accuracy of 98% and pair-wise average Cohen’s Kappa () score of 0.65 (substantial agreement), showcasing the high reliability of the assigned output label. Table 4 shows the average score over all the designed aspects to measure the reliability of the proposed dataset. Human evaluation across various designed aspects consistently yields an average score greater than 4.5 out of 5, which underscores the reliability and quality of the EngSAF dataset. The details of the human annotators can be found in Technical Appendix.
4 Experiments
Building on the foundational work by Filighera et al. (2022) in their investigation, our primary objective is to conduct experiments to establish a baseline for our EngSAF dataset. We also seek to delve deeper into the impact of incorporating questions on the generated feedback and the assigned labels. Traditionally, in Automated Short Answer Grading (ASAG), assessments have focused solely on evaluating reference answers and student responses. However, Lv et al. (2021) challenge this convention by demonstrating that integrating questions into the evaluation process enhances the performance of traditional ASAG tasks. Therefore, our study aims to build upon these insights to explore the nuanced effects of incorporating questions in ASAG assessments, aiming to refine and improve the evaluation criteria and methodologies in this domain.
4.1 Experimental Setting
To establish baselines for EngSAF, we have conducted experiments in two settings.
1.
Fine-tuning Large Language Models (LLMs): We use the Llama-2 model Touvron et al. (2023), Mistral 7b model555https://mistral.ai/, which are fine-tuned to predict the output label for student responses, categorizing them as correct, partially correct, or incorrect and jointly providing feedback explaining the assigned output label. Furthermore, we have conducted this experiment using two distinct methodologies.
Without Question: Student answer and Reference answer are passed as input.
With Question: Question, Student answer, and Reference answer are passed as input.
The following models are used in this experiment.
llama-2-13B-chat
Mistral-7B-Instruct-v0.1 -
2.
Zero-shot experiment using chatGPT: For this experiment, we prompted ChatGPT to assign an output label along with feedback to each student’s response by evaluating its correctness compared to the reference answer for the ASAG task. Prompt details are present in the Techinical Appendix. We used ChatGPT API666https://openai.com/ to perform this experiment on EngSAF Dataset. The following model has been used in this experiment.
gpt-3.5-turbo-16k
Hyperparameters used for fine-tuning LLMs on EngSAF can be found in Technical Appendix 10.2. All experiments were performed on 2 NVIDIA A100-SXM GPU with 80 GiB of memory. Fine-tuning takes about 1 hour per epoch for training.
5 Results
Table 5 shows a majority baseline, Llama-2 fine-tuning, Mistral-7b fine-tuning, and ChatGPT zero-shot experiments results. The majority baseline contains the most occurring label and feedback from the EngSAF train set. The most common label is “correct response,” and the most common feedback is “Well done! You have answered the question correctly, covering all the required aspects.”. We conducted the ChatGPT experiment only on the unseen answers test set due to its zero-shot setting, as both unseen questions and answers are identical in the zero-shot setting.
| Unseen Answers | Unseen Questions | |||||||||||
| Model | Acc. | F1 | BLEU | MET. | ROU. | BERT | Acc. | F1 | BLEU | MET. | ROU. | BERT |
| Majority | 43.3 | 26.1 | 1.2 | 8.6 | 12.7 | 16.2 | 40.5 | 23.4 | 0.1 | 8.64 | 2.78 | 12.32 |
| 74.4 | 73.0 | 13.3 | 34.1 | 16.8 | 31.9 | 55.6 | 53.6 | 12.5 | 31.4 | 16.2 | 28.9 | |
| 73.9 | 73.7 | 11.7 | 35.9 | 16.9 | 35.1 | 56.3 | 54.9 | 9.2 | 31.4 | 13.9 | 32.6 | |
| 72.8 | 73.1 | 11.3 | 33.3 | 14.82 | 37.4 | 54.7 | 55.4 | 11.3 | 32.0 | 15.2 | 36.2 | |
| 75.4 | 75.7 | 13.9 | 38.3 | 19.5 | 41.6 | 58.7 | 57.9 | 14.9 | 36.7 | 19.7 | 40.9 | |
| ChatGPT | 41.1 | 24.0 | 7.4 | 31 | 11.3 | 28.8 | - | - | - | - | - | - |
5.1 Analysis
In this section, we delve into the insights drawn from the performance results of experiments on the EngSAF dataset for the ASAG task. The evaluation metrics used for labels are accuracy and macro-averaged F1. To evaluate the feedback, we measure the Rouge-2 Post (2018), SCAREBLEU 777https://pypi.org/project/sacrebleu/, METEOR Banerjee and Lavie (2005) and BERTSCORE Zhang et al. (2019) scores.
Table 5 shows that Mistral-7b significantly outperforms the majority baseline in both label and feedback metrics. However, a notable disparity exists in improvement between unseen questions and unseen answers. We observe a decrease of approximately 23% in accuracy for unseen questions compared to unseen answers. This underscores the challenge even fine-tuned models face when tasked with generalizing to new questions. Additionally, there is approximately 4% decrease in Bertscore for unseen questions compared to unseen answers, suggesting the necessity for new evaluation metrics that comprehensively assess text on the context level instead of the lexical level. Including questions in the input has led to a notable improvement in Bertscore, with a 10% increase for unseen answers and a 12% increase for unseen questions. Similarly, for label accuracy, we observe a 3.5% increase for unseen answers and a 7.3% increase for unseen questions in the test set. Technical Appendix includes qualitative analysis of test examples. Overall, the experiment suggests including questions for improvement in the predicted feedback by the model. The ChatGPT zero-shot experiment yields the lowest accuracy at 41.1%, even falling below that of the majority baseline, highlighting the complexity of the ASAG task and struggle of large language models like ChatGPT struggle in a zero-shot setting. Another contributing factor to this performance disparity could be the inherent bias present in teacher grading, which models can only learn when trained on such data. Interestingly, the smaller Mistral 7B model outperformed the LLaMA-2-13B model for ASAG, demonstrating the potential of more compact models in this task.
6 Deployment
| Aspect | Avg. Score |
| Output Label Accuracy | 92.5% |
| Feedback Quality/Correctness | 4.5 |
| Emotional Impact | 4.9 |
For the real-world deployment, our fine-tuned ASAG model was integrated into an end-semester exam on ET 623 (Learning Analytics Course) at IIT Bombay for the 2024 academic year. The setup included the students enrolled in the course who consented to participate in the experiment. The end-sem exam includes 2 short-answer questions, each accompanied by the instructor’s correct/reference answer. We randomly sampled 25 student answers for each question and used our fine-tuned ASAG with Feedback model to predict the output label and feedback/ explanation for the predicted output. Each predicted output label was evaluated based on its correctness, where the subject matter expert checks whether the predicted output label is correct or not. Each predicted feedback is analyzed in terms of Feedback Correctness/ Accuracy and Emotional Impact as discussed in the Quality Estimate section 3.5. Each aspect is scored on a scale (1-5), with a high score indicating a better response.
Table 6 presents the average scores across all the designed aspects used to measure the reliability of the proposed dataset. Upon examining the table, it is evident that the subject matter evaluation scores greater than 4.5 for both Feedback Quality/Correctness and the Emotional Impact aspect. This demonstrates the reliability and effectiveness of the ASAG model in real-world scenarios. Additionally, the model’s predicted output label achieved an accuracy of 92.5%, further showcasing its performance and reliability. Technical Appendix contains more details about deployment.
7 Summary, Conclusion, and Future Work
This study presents a novel and extensive dataset for grading short answers across multiple engineering domains. Our dataset addresses the need for a standardized evaluation platform by covering diverse subjects and a wide range of short answers. We meticulously curated and annotated this dataset to ensure its quality and applicability. Using the Label Aware Synthetic Feedback Generation (LASFG) strategy we have synthetically incorporated feedback in the EngSAF dataset. Through experimentation, we benchmarked the dataset’s performance using different LLM’s fine-tuning and chatGPT zero shot experiments.
In conclusion, our research endeavors culminated in successfully creating and evaluating a cutting-edge multi-domain short answer grading dataset. The dataset’s diverse content and meticulous annotations provide a solid platform for training and assessing grading models across various subject domains. Our benchmarking experiments showcased promising results, indicating that our dataset has the potential to significantly improve the quality and reliability of ASAG systems. Additionally, we deployed the ASAG system in a real-world setting for an IITB course exam, demonstrating its practical applicability. This work marks a crucial step forward in educational technology, equipping educators and researchers with a valuable resource to boost advancements in automated assessment techniques.
While this research presents a novel dataset for effective short answer grading along with feedback, there are several possibilities for future investigation. One area of focus could be the refinement and expansion of our dataset to include more nuanced and complex short answers, enabling us to accommodate a broader range of grading scenarios. Baseline figures could be improved by expansion or augmentation of the dataset. External knowledge inclusion through knowledge graphs could be explored in this direction.
8 Limitations
Our study, titled “I understand why I got this grade": Automatic Short Answer Grading with Feedback" has provided valuable insights into the multi-domain ASAG. However, it’s important to acknowledge certain limitations. The synthetic nature of the generated feedback may introduce biases or might contain inconsistencies, factual inaccuracies, or overly generic content, that could impact the implications of the results. One key finding from table 3 is an imbalance in output labels. Specifically, Output label Zero (0) is less frequent compared to labels One (1) and Two (2). This imbalance could impact the performance of automatic grading models trained on this dataset. Despite our efforts to curate diverse short-answer responses across multiple engineering domains, the dataset’s overall size might limit the complexity and depth of models that can be trained on it. Variations in data distribution, language patterns, and task requirements across different domains and contexts may impact the application of our results to real-world scenarios. While our dataset’s strength lies in its coverage of multiple engineering domains, it’s important to note that the range of domains might not be exhaustive. Researchers focusing on specific domains or aiming to enhance domain-specific performance may need to consider domain-specific or subject-specific fine-tuning or gather additional data.
9 Ethical Statement
The EngSAF dataset contains questions, reference answers, and student answers along with the output label assigned for each student answer. This dataset is taken from actual exams and quizzes conducted at the university. We are committed to protecting those student’s privacy and anonymity. Hence, no student-specific information is disclosed within the dataset. The process of collecting and curating this dataset adhered to the ethical norms of informed consent. Prior to any contribution, instructors of each course were provided details concerning the data collection’s intent, the utilization of their questions and responses, and any potential implications. While the ASAG models created with this dataset offer great potential in educational assessment, they must be used properly and ethically. As with any AI technology, its impact on education must be regularly monitored and evaluated. Ensuring an appropriate balance between the advantages derived from automation and safeguarding the educational experience and human judgement is of utmost importance.

### v2

@@
-"I understand why I got this grade": Automatic Short Answer Grading with Feedback
+“I understand why I got this grade”: Automatic Short Answer Grading (ASAG) with Feedback
 Abstract
-The demand for efficient and accurate assessment methods has intensified as education systems transition to digital platforms.
-Providing feedback is essential in educational settings and goes beyond simply conveying marks as it justifies the assigned marks.
-In this context, we present a significant advancement in automated grading by introducing Engineering Short Answer Feedback (EngSAF)—a dataset of student answers accompanied by reference answers and questions for the Automatic Short Answer Grading (ASAG) task.
-The EngSAF dataset is meticulously curated to cover a diverse range of subjects, questions, and answer patterns from multiple engineering domains.
-We leverage state-of-the-art large language models’ (LLMs) generative capabilities with our Label-Aware Synthetic Feedback Generation (LASFG) strategy to include feedback in our dataset.
-This paper underscores the importance of enhanced feedback in practical educational settings, outlines dataset annotation and feedback generation processes, conducts a thorough EngSAF analysis, and provides different LLMs-based zero-shot and finetuned baselines for future comparison111Our code and dataset are available at https://github.com/dishankaggarwal/EngSAF.
-Additionally, we demonstrate the efficiency and effectiveness of the ASAG system through its deployment in a real-world end-semester exam at the Indian Institute of Technology Bombay (IITB), showcasing its practical viability and potential for broader implementation in educational institutions.
-1 Introduction
-Technology integration in education has resulted in transformative changes, redefining traditional pedagogical approaches and assessment methodologies.
-Effective education relies on feedback and explanations provided during assessments to ensure quality learning outcomes Shute (2008).
-Grading questions in tests and examinations have proven to be a good measure to assess student’s learning and understanding of a topic or a subject.
-An exam could include various question types, such as multiple choice, fill-in-the-blank, short answers, essays, etc.
-Among these question types, short answers and essays are more complicated to analyze than multiple-choice or fill-in-the-blank type questions due to flexibility and natural language in the response.
-Automating the grading process becomes crucial, especially in countries with extremely high student-to-teacher ratios, as it can significantly reduce instructor’s workloads and improve the assessment process.
-Significant advancements have been achieved in this field in recent years, primarily after the introduction of SemEval Semantic Textual Similarity (STS) task Agirre et al.
-(2012, 2013, 2014, 2015).
-This challenge can be approached as a machine learning issue, where the objective is to grade a student’s response based on how similar it is to the reference answers.
-However, simply assigning a score or label to a learner’s response is often inadequate in practical educational contexts.
+In recent years, there has been a growing interest in using Artificial Intelligence (AI) to automate student assessment in education.
+Among different types of assessments, summative assessments play a crucial role in evaluating a student’s understanding level of a course.
+Such examinations often involve short-answer questions.
+However, grading these responses and providing meaningful feedback manually at scale is both time-consuming and labor-intensive.
+Feedback is particularly important, as it helps students recognize their strengths and areas for improvement.
+Despite the importance of this task, there is a significant lack of publicly available datasets that support automatic short-answer grading with feedback generation.
+To address this gap, we introduce Engineering Short Answer Feedback (EngSAF), a dataset designed for automatic short-answer grading with feedback.
+The dataset covers a diverse range of subjects, questions, and answer patterns from multiple engineering domains and contains data points.
+We incorporate feedback into our dataset by leveraging the generative capabilities of state-of-the-art large language models (LLMs) using our Label-Aware Synthetic Feedback Generation (LASFG) strategy.
+This paper underscores the importance of enhanced feedback in practical educational settings, outlines dataset annotation and feedback generation processes, conducts a thorough EngSAF analysis, and provides different LLMs-based zero-shot and finetuned baselines for future comparison111Dataset and code are open-sourced at: https://github.com/dishankaggarwal/EngSAF.
+The best-performing model (Mistral-7B) achieves an overall accuracy of 75.4% and 58.7% on unseen answers and unseen question test sets, respectively.
+Additionally, we demonstrate the efficiency and effectiveness of our ASAG system through its deployment in a real-world end-semester exam at a reputed institute.
+There, we achieve an output label accuracy of 92.5% along with feedback correctness and emotional impact scores above 4.5 (out of 5) on human evaluation, thus showcasing its practical viability and potential for broader implementation in educational institutions.
+Keywords:
+LLM Automatic Short Answer Grading EngSAF Feedback1 Introduction
+Technology integration in education, particularly artificial intelligence (AI), has resulted in transformative changes that have redefined traditional pedagogical approaches and assessment methodologies.
+Effective education relies on feedback and explanations provided during assessments to ensure quality learning outcomes [30].
+In the context of education, educators use different kinds of assessments, including formative, summative, and diagnostic222https://www.niu.edu/citl/resources/guides/instructional-guide/formative-and-summative-assessment.shtml, each of which serves a vital role in a student’s learning journey.
+This work deals with summative assessments, which are essential in determining whether the student has achieved the learning goals of a course once it is over.
+Grading summative assessments often involve grading short answers and essays, which are more complicated in nature due to the flexibility and natural language in the response.
+Hence, automating the grading process for such questions becomes crucial, especially in countries with extremely high student-to-teacher ratios, as it can significantly reduce instructor’s workloads and improve the assessment process.
+For instance, during the COVID-19 pandemic, educators faced increased stress and exhaustion due to overloaded grading burden [15, 25].
+This challenge can be approached as a machine learning problem, where the objective is to grade a student’s response based on how similar it is to the reference answers.
+However, simply assigning a score or label to a learner’s response is often inadequate in practical educational contexts [13, 31].
 Ahea et al.
-(2016) highlights the value and effectiveness of feedback in improving students’ learning and professionalizing teaching in higher education.
-1.1 Problem Statement
-Given a question, a reference answer, and a student’s answer, the aim is to provide content-focused elaborated feedback and assign a label indicating the degree of correctness in the student’s answer compared to the reference answer.
-Here, we focus on questions where the answer type is a sentence or a short paragraph.
-This task involves evaluating the alignment between the student’s and reference answers.
+[5] highlights the value and effectiveness of feedback in improving student’s learning and standardizing teaching in higher education.
+Although earlier work formulated this problem as the Semantic Textual Similarity (STS) task [3, 4, 2, 1], recent works acknowledge the importance of feedback generation along with the grading task [12].
+Hence, together with the grading task and the feedback generation task, researchers have formulated the problem of automatic short answer grading (ASAG) with feedback.
+1.1 Automatic Short Answer Grading (ASAG) with Feedback
+The ASAG with feedback problem is as follows: given a question, a reference answer, and a student’s answer, the aim is to assign a label indicating the degree of correctness in the student’s answer compared to the reference answer and provide content-focused elaborated feedback/explanation for the same.
+Note that the degree of correctness is limited to only three labels here, namely correct, partially correct, and incorrect.
+In this problem, we only focus on questions where the answers are short, varying between a sentence and a short paragraph.
+This task involves evaluating the alignment between the student’s answer and the reference answers which is expressed via the degree of correctness along with proper reasoning as to why that label is assigned.
 Figure 1 illustrates the problem statement, showing the input and output using an example from the EngSAF dataset.
+1.2 Motivation
+With recent advancements in state-of-the-art (SOTA) Natural Language Processing (NLP) techniques, there is an increasing interest in applying such methodologies in the field of education, particularly in short-answer evaluations.
+The benefits of designing such AI-assisted systems that solve the ASAG with feedback problem are two-fold:
+1.
+The learners333We often use the terms student and learner synonymously.
+Also, we do the same for the terms graders, teachers, and educators.
+benefit by understanding their strengths and shortcomings from the content-elaborated feedback provided by the system.
+2.
+The educators benefit by having prior feedback on the student’s degree of correctness, along with a proper explanation for the same.
+This can reduce an educator’s cognitive load while grading such short answer questions.
+However, researchers face challenges when designing such systems due to the lack of public, content-centered, elaborated feedback datasets in different domains.
+These datasets are crucial for training and developing automated grading systems that can provide feedback.
+Thus, this calls for a more efficient and effective dataset spanning multiple engineering domains along with proper grades and content-elaborated feedback.
+This is where we introduce the Engineering Short Answer Feedback (EngSAF) dataset.
+Our contributions are:
+1.
+A Label Aware Synthetic Feedback Generation (LASFG) strategy to convert conventional automatic short answer grading datasets to the ones containing feedback. (Section 3.2)
+2.
+EngSAF dataset containing around 5.8K student responses to 119 questions from multiple engineering domains taken from real-life examinations at a reputed institute along with synthetically generated feedback using LASFG strategy explaining the assigned grade for the task of ASAG.
+To the best of our knowledge, this is the first dataset containing questions and responses from multiple engineering domains.
+The overarching goal of this work is to provide helpful feedback to students automatically.
+(Section 3)
+3.
+Benchmark scores on the EngSAF dataset using different Large Language Models (LLMs) for future comparison and research. (Section 5)
+4.
+Real-world deployment of the best-performing fine-tuned ASAG model in an end-semester exam at a reputed institute. (Section 6)
+2 Related work
+This section provides a comprehensive literature review, beginning with previous research on the ASAG problem, followed by an in-depth survey of existing short-answer grading datasets.
+2.1 Automatic Short Answer Grading (ASAG)
+ASAG is an essential area of research that has garnered significant attention in recent years.
+Several approaches have been proposed for traditional ASAG techniques, ranging from rule-based methods to more sophisticated machine-learning techniques.
+One early approach for solving the ASAG problem was based on keyword or pattern matching, where the presence or absence of certain keywords in the student’s answer was used to determine its accuracy [20, 32, 23].
+To overcome these limitations, researchers have developed more sophisticated methods that use NLP-based techniques.
+One such method is based on Latent Semantic Analysis (LSA), which represents texts as high-dimensional vectors and compares them to the reference answers using cosine similarity [17].
+In a related study, the task of ASAG is addressed by incorporating features such as answer length, grammatical correctness, and semantic similarity in comparison to reference answers [33].
+More recently, deep learning models such as convolutional neural networks (CNNs) and recurrent neural networks (RNNs) have been applied to ASAG’s task.
+These models are trained on large amounts of annotated data and can capture the semantic relationships between words in a student’s answer and the reference answers [34, 36].
+Pre-trained Language Models (PLMs), such as BERT [8], GPT [27], RoBERTa [18], DistillBERT [29], ALBERT [16], and SBERT[28] have performed exceptionally well for various NLP tasks, are widely used even for the task of ASAG on traditional ASAG datasets [7].
+A more recent work by Fateen et al.
+[11] extends this problem by generating feedback while solving the ASAG problem using a RAG-based approach.
+However, a common limitation of all the prior works is that only grades are assigned, and none of the datasets provide a ground truth for the problem content-elaborated feedback generation.
+2.2 Short Answer Grading Datasets
+Again, when it comes to short answer grading datasets, researchers have carefully curated several publicly accessible datasets to facilitate research and benchmarking of different solutions on the ASAG task.
+In 2009, Mohler et al.
+[22] published a short answer dataset on a data structures course that contained 630 records.
+In 2011, the same authors [21] published an extended dataset on the same course that contained 2273 records and named it the Texas Extended dataset.
+Apart from this, Nielsen et al.
+[24] created the Beetle corpus, which contains fine-grained annotations of entailment relationships in the context of student answers to science assessment questions.
+Another dataset that is publicly available on Kaggle and could be used for the task of ASAG is provided by Hewlett Foundation named ASAP-AES444https://www.kaggle.com/c/asap-aes/data (Automated Assessment Prize Competition for Essay Scoring).
+However, as mentioned above, all datasets are conventional ASAG datasets that solely consist of grades or scores.
+Our extensive literature survey includes only one short-answer grading dataset with content-focused elaborated feedback, i.e., [12], which introduces an inaugural dataset for this problem and comprises bilingual responses in English and German.
+Human annotations were meticulously collected and refined to uphold the quality of the feedback.
+However, this dataset is constrained by its limited number of student responses () in English and its exclusive focus on questions from a single domain, specifically computer network queries, thereby lacking diversity across different engineering domains.
+To address these limitations, we curate the EngSAF dataset, which contains data points and questions from multiple engineering domains.
+3 Engineering Short Answer Feedback (EngSAF) Dataset
+This dataset contains 119 questions drawn from different undergraduate and graduate engineering courses, accompanied by approximately 5.8k student responses.
+An instructor provided correct/reference answer also accompanies each question.
+These questions and answers have been taken from real-life examinations at a reputed institute, thus covering a diverse range of 25 courses.
+The questions and responses span across multiple domains, including image processing, water quality management, and operating systems, to name a few.
+The dataset has been curated using a well-known e-learning platform that has been extensively used and developed at a reputed institute.
+Based on the instructor’s assigned marks, each response has been categorized as “correct,” “partially correct,” or “incorrect.”
 | Question | What is the difference between basin order and channel order? |
@@
 |
-1.2 Motivation
-The increasing demand for technology in education has led to a need for more efficient and effective methods of grading short-answer assessments.
-In the case of short answer grading, feedback with clear explanations goes beyond simply conveying a grade.
-It offers valuable insights into student’s strengths, weaknesses, and areas of improvement.
-However, the effectiveness of feedback is largely unexplored due to the lack of public, content-centered, elaborated feedback datasets in different domains.
-These datasets are crucial for training and developing automated feedback systems that can provide personalized and nuanced feedback.
-The need for a more efficient and effective ASAG system incorporating feedback alongside grade has arisen due to these limitations, and this is where the Engineering Short Answer Feedback (EngSAF) dataset steps in, consisting of questions students answer from multiple engineering domains for the ASAG task.
-Our Contributions are:
-1.
-EngSAF dataset containing around 5.8K student responses to 119 questions from multiple engineering domains along with synthetically generated feedback explaining the assigned grade for the task of ASAG.
-To the best of our knowledge, this is the first dataset containing questions and responses from multiple engineering domains.
-(Section 3)
-2.
-Benchmark scores on the EngSAF dataset using different Large Language Models (LLM) for future comparison and research. (Section 5)
-3.
-Real-world deployment of the EngSAF fine-tuned ASAG model in an end-semester exam at Indian Institute of Technology, Bombay (IITB) (Section 6).
-2 Related work
-2.1 Automatic Short Answer Grading (ASAG)
-ASAG is an essential area of research that has garnered significant attention in recent years.
-Several approaches have been proposed for traditional ASAG, ranging from rule-based methods to more sophisticated machine-learning techniques.
-One early approach for ASAG was based on keyword or pattern matching, where the presence or absence of certain keywords in the student’s answer was used to determine its accuracy.
-Mitchell et al.
-(2002); Sukkarieh et al.
-(2004); Nielsen et al.
-(2009).To overcome these limitations, researchers have developed more sophisticated methods that use natural language processing (NLP) techniques.
-One such method is based on Latent Semantic Analysis (LSA), which represents texts as high-dimensional vectors and compares them to the reference answers using cosine similarity LaVoie et al.
-(2020).
-In a related study, the task of ASAG is addressed by incorporating features such as answer length, grammatical correctness, and semantic similarity in comparison to reference answers Sultan et al.
-(2016).
-More recently, deep learning models such as convolutional neural networks (CNNs) and recurrent neural networks (RNNs) have been applied to ASAG’s task.
-These models are trained on large amounts of annotated data and can capture the semantic relationships between words in a student’s answer, and the reference answers Surya et al.
-(2019); Zhang et al.
-(2022).
-Pre-trained Language Models (PLMs), such as BERT Devlin et al.
-(2018), GPT Radford et al.
-(2019), RoBERTa Liu et al.
-(2019), DistillBERT Sanh et al.
-(2019), and ALBERT Lan et al.
-(2019) have performed exceptionally well for various tasks in NLP, also used for the task of ASAG.
-Sentence-BERT or SBERT Reimers and Gurevych (2019) performs exceptionally well on current publicly available traditional ASAG datasets Condor et al.
-(2021).
-A common limitation of the prior work is that only grades are assigned, and no feedback is provided.
-We seek to fill this gap with the introduction of the EngSAF dataset.
-2.2 Short Answer Grading Datasets
-Several publicly accessible datasets have been curated, each designed to facilitate research and benchmarking in the ASAG task.
-In 2009, Mohler and Milhalcea published a short answer dataset on a data structures course that contained 630 records Mohler and Mihalcea (2009).
-In 2011, they published an extended dataset on the same course that contained 2273 records Mohler et al.
-(2011), named the Texas Extended dataset.
-To complete the assignment of student-authored answer annotation, the Beetle corpus is being created Nielsen et al.
-(2008).
-Another dataset that is publicly available on Kaggle and could be used for the task of ASAG is provided by Hewlett Foundation named ASAP-AES222https://www.kaggle.com/c/asap-aes/data (Automated Assessment Prize Competition for Essay Scoring).
-However, all aforementioned datasets are conventional ASAG datasets solely consisting of grades or scores.
-Our extensive literature survey includes only one short answer grading dataset with content-focused elaborated feedback, i.e., Filighera et al.
-(2022), which introduces an inaugural dataset for short-answer feedback comprising bilingual responses in English and German.
-Human annotations were meticulously collected and refined to uphold the feedback’s quality.
-However, this dataset is constrained by its limited number of student responses () in English and its exclusive focus on questions from a single domain, specifically computer network queries, thereby lacking diversity across different engineering domains.
-To address these limitations, we introduced EngSAF, which contains data points and questions from multiple engineering domains.
-3 Engineering Short Answer Feedback (EngSAF) Dataset
-This dataset contains 119 questions drawn from different undergraduate and graduate engineering courses, accompanied by approximately 5.8k student responses.
-An Instructor provided correct answer/ reference answer also accompanies each question.
-These questions and answers have been taken from actual quizzes/ exams from the Indian Institute of Technology, Bombay (IITB), covering a diverse range of 25 courses.
-The questions and responses span across multiple domains, including image processing, water quality management, and operating systems, to name a few.
-Technical Appendix contains the complete list of courses included in the dataset.
-Based on the instructor’s assigned marks, each response has been categorized as “correct,” “partially correct,” or “incorrect.”
-Table 1 provides an example from the EngSAF dataset showcasing a question, its corresponding reference answer, and three different student responses alongside their associated output labels and synthetically generated feedback/Explanation.
-Technical Appendix contains more examples from the EngSAF dataset.
+Table 1 provides examples from the EngSAF dataset showcasing a question, its corresponding reference answer, and three different student responses alongside their associated output labels and synthetically generated feedback.
 The dataset is applicable for both traditional automatic short answer grading and the generation of elaborated feedback.
 3.1 Dataset Construction
-EngSAF contains questions and student responses from different undergraduate and graduate engineering courses.
-The instruction provided a correct answer/ reference answer accompanying each question.
-The Instructor/ Teaching Assistant (TA) assigns marks to each student’s answer from their respective course.
-The students’ responses and grades are obtained from a reputed engineering university.
+The EngSAF dataset contains questions and student responses from different undergraduate and graduate engineering courses from a reputed institute.
+The data was collected using a well-known e-learning platform that has been developed and is extensively used at this institute.
+Using the e-learning platform, the instructor provided a correct answer/ reference answer to each question.
+Additionally, the instructors and teaching assistants (TAs) have assigned marks to each student’s answer from their respective courses.
+The student’s responses and grades were then obtained from the e-learning platform.
 The quality assessment of the proposed dataset is covered in subsection 3.4.
-If the maximum mark for a question is and a student’s answer is graded marks by the instructor, then the output label is annotated according to the scheme in Table 2.
 | Condition | Output Label |
@@
 | 0 (Incorrect Response) |
-3.2 Challenges and Requirements
-Diversity in student responses: Human language is inherently subjective, and different individuals may interpret the same content differently.
-This subjectivity extends to the grading process, making it challenging to provide universally applicable feedback.
-Response Variability: Short answer responses can vary significantly regarding quality, coherence, and relevance.
-Some responses may be well-structured and articulate, while others may be incomplete or contain grammatical errors.
-Designing feedback that caters to this variability in response quality poses a considerable challenge.
-Feedback Impact: Wrong feedback can negatively shape learners’ perceptions of themselves and their abilities.
-If the feedback consistently highlights their shortcomings or mistakes, they may internalize these negative perceptions and develop a fixed mindset about their capabilities.
-This can hinder their willingness to take risks, seek challenges, and persist in facing difficulties.
+Each data point includes a question, student answer, and reference answer, along with a label indicating the level of correctness of the student’s response and detailed, content-specific feedback.
+If the maximum mark for a question is and a student’s answer is graded marks by the instructor, then the output label is annotated according to the scheme in Table 2.
 3.2 Label Aware Synthetic Feedback Generation (LASFG)
-Leveraging the advanced language generative and reasoning capabilities of state-of-the-art Large Language Models (LLMs) like Gemini333https://gemini.google.com/ and ChatGPT, we enhance educational assessment with feedback, particularly in Automatic Short Answer Grading (ASAG).
+Apart from the question, the student’s answer, and the reference answer, each data point contains a content-elaborated feedback part.
+To obtain this, we leverage the advanced language generative and reasoning capabilities of SOTA LLMs like Gemini555https://gemini.google.com/ and ChatGPT.
 The approach involves utilizing Gemini’s ability to comprehend input prompts consisting of a question, a student’s answer, a reference answer, and the corresponding grading label provided by an instructor or teaching assistant to generate content-focused elaborated feedback.
 The generated feedback covers the reasoning or explanation of the Gold output label.
-Grammarly444https://www.grammarly.com/ then checks the synthetically generated feedback to remove any grammatical errors.
-Prompt details used for synthetic feedback generation can be found in the Technical Appendix 10.1.
-The quality estimation of the generated feedback is shown in Section 3.5
+Based on initial observations of the generated feedback quality, we use gemini-1.5-flash model for feedback generation.
+To ensure the quality of the generated feedback, we performed quality estimation as described in Section 3.4.
 3.3 Corpus Statistics
+Following Dzikovska et al.’s approach [10], we partitioned the data into training sets, comprising 70% of the dataset, as well as unseen answers (16%) and unseen questions (14%) for test sets as shown in Table 3.
+The test split with unseen answers (UA) includes fresh responses to the questions used in training.
+In contrast, the test split with unseen questions (UQ) comprises entirely new questions for testing the model’s ability to generalize to new questions without prior exposure.
+The average text length (in tokens) for different fields in our dataset is as follows: Questions have an average length of 17.50 tokens, student answers average 25.85 tokens, reference answers are 26.47 tokens long on average, and feedback is the longest, with an average of 41.95 tokens.
 | Label | Train | UA | UQ | Total |
@@
 | Total | 4069 | 980 | 765 | 5814 |
-Following Dzikovska et al.
-(2013), we partitioned the data into training sets, comprising 70% of the dataset, as well as unseen answers (16%) and unseen questions (14%) for test sets as shown in table 3.
-The test split with unseen answers includes fresh responses to the questions used in training.
-In contrast, the test split with unseen questions comprises entirely new questions for testing the model’s ability to generalize to new questions without prior exposure.
-Technical Appendix 10.3 contains insight about insights about the distribution of text length for student answers, reference answers, questions, and feedback in the EngSAF dataset.
 3.4 Quality Estimation
@@
 This sampled data is a good representation of the EngSAF dataset.
-Each synthetically generated feedback was scored by three human annotators over three aspects.
-Each aspect is scored on a scale (1-5), with a high score indicating a better response.
-Each feedback is analyzed on the following aspects.
+Each synthetically generated feedback was evaluated on three aspects by three proficient English-speaking annotators, all Master’s students.
+They were provided with detailed annotation guidelines and examples to ensure consistency.
+Each aspect is scored on a scale (1-5), with a high score indicating a better response. Each feedback is analyzed on the following aspects.
 1.
-Fluency and Grammatical Correctness: This aspect tests whether the generated feedback is fluent in English and grammatically correct.
+Fluency and Grammatical Correctness (FGC): This aspect checks if the feedback is fluent and grammatically correct in English.
 2.
-Feedback Correctness/Accuracy: This aspect assesses the overall quality of the generated feedback regarding content, relevance, quality, and explanation for the assigned grade.
+Feedback Correctness/Accuracy (FC): This aspect assesses the overall quality of the generated feedback regarding content, relevance, quality, and explanation for the assigned grade.
 3.
-Emotional Impact: This aspect assesses how feedback affects the learner’s emotional state.
-Annotators are tasked with rating whether the feedback avoids triggering negative emotions or impacts by refraining from using words such as "fail" that may evoke feelings of discouragement or distress in the learner.
+Emotional Impact (EI): This aspect assesses how feedback affects the learner’s emotional state.
+Annotators are tasked with rating whether the feedback avoids triggering negative emotions or impacts by refraining from using words such as “fail” that may evoke feelings of discouragement or distress in the learner.
 | Aspect | Avg. Score |
-| Fluency & Grammatical Correctness | 4.73 |
-| Feedback Correctness/Accuracy | 4.55 |
-| Emotional Impact | 4.61 |
+| Fluency & Grammatical Correctness (FGC) | 4.73 |
+| Feedback Correctness/Accuracy (FC) | 4.55 |
+| Emotional Impact (EI) | 4.61 |
 Three human annotators evaluate the correctness of each output label for sampled data points, achieving an accuracy of 98% and pair-wise average Cohen’s Kappa () score of 0.65 (substantial agreement), showcasing the high reliability of the assigned output label.
@@
 Human evaluation across various designed aspects consistently yields an average score greater than 4.5 out of 5, which underscores the reliability and quality of the EngSAF dataset.
-The details of the human annotators can be found in Technical Appendix.
+As part of our ablation studies, we have also performed an LLM-based evaluation by following the LLM-as-a-judge paradigm [38].
+We conducted such an evaluation using the Gemini model to assess the quality of the generated feedback.
+Specifically, we instructed Gemini to evaluate feedback quality on the EngSAF test set, and we obtained average scores of 4.23 for feedback quality on unseen answers (UA) and 3.91 on unseen questions (UQ) out of 5.
+These scores reflect the model’s assessment of the feedback quality and provide an additional layer of validation to the reliability and quality of the dataset.
 4 Experiments
-Building on the foundational work by Filighera et al.
-(2022) in their investigation, our primary objective is to conduct experiments to establish a baseline for our EngSAF dataset.
-We also seek to delve deeper into the impact of incorporating questions on the generated feedback and the assigned labels.
-Traditionally, in Automated Short Answer Grading (ASAG), assessments have focused solely on evaluating reference answers and student responses.
+Following Filighera et at.’s [12] methodology, our primary objective is to establish a baseline for our EngSAF dataset.
+We also assess the impact of incorporating questions on the generated feedback and the assigned labels, by performing experiments in two different settings, namely with_question and without_question.
+Traditionally, in Automatic Short Answer Grading (ASAG), assessments have focused solely on evaluating reference answers and student responses.
 However, Lv et al.
-(2021) challenge this convention by demonstrating that integrating questions into the evaluation process enhances the performance of traditional ASAG tasks.
-Therefore, our study aims to build upon these insights to explore the nuanced effects of incorporating questions in ASAG assessments, aiming to refine and improve the evaluation criteria and methodologies in this domain.
+[19] challenge this convention by demonstrating that integrating questions into the evaluation process improves the performance of traditional ASAG tasks.
 4.1 Experimental Setting
@@
 1.
-Fine-tuning Large Language Models (LLMs): We use the Llama-2 model Touvron et al.
-(2023), Mistral 7b model555https://mistral.ai/, which are fine-tuned to predict the output label for student responses, categorizing them as correct, partially correct, or incorrect and jointly providing feedback explaining the assigned output label.
+Fine-tuning Large Language Models (LLMs): We use the Llama-2 model [35], Llama-3.1 model [9], Mistral 7b model666https://mistral.ai/, which are fine-tuned to predict the output label for student responses, categorizing them as correct, partially correct, or incorrect and jointly providing feedback explaining the assigned output label.
 Furthermore, we have conducted this experiment using two distinct methodologies.
@@
 With Question: Question, Student answer, and Reference answer are passed as input.
-The following models are used in this experiment.
-llama-2-13B-chat
-Mistral-7B-Instruct-v0.1 -
+The models used in this experiment are llama-3.1-8B-Instruct-chat, llama-2-13B and Mistral-7B-Instruct-v0.1.
 2.
-Zero-shot experiment using chatGPT: For this experiment, we prompted ChatGPT to assign an output label along with feedback to each student’s response by evaluating its correctness compared to the reference answer for the ASAG task.
-Prompt details are present in the Techinical Appendix.
-We used ChatGPT API666https://openai.com/ to perform this experiment on EngSAF Dataset.
-The following model has been used in this experiment.
-gpt-3.5-turbo-16k
-Hyperparameters used for fine-tuning LLMs on EngSAF can be found in Technical Appendix 10.2.
-All experiments were performed on 2 NVIDIA A100-SXM GPU with 80 GiB of memory.
-Fine-tuning takes about 1 hour per epoch for training.
+Zero-shot experimentation: For this experiment, we prompt ChatGPT777https://openai.com/ and DeepSeek [14] to assign an output label along with feedback to each student’s response by evaluating its correctness compared to the reference answer for the ASAG task.
+The following models are used: gpt-4o and DeepSeek-R1
+All experiments are performed on 2 NVIDIA A100-SXM GPU with 80 GB of memory. Fine-tuning takes around 2 hours per epoch for training.
 5 Results
-Table 5 shows a majority baseline, Llama-2 fine-tuning, Mistral-7b fine-tuning, and ChatGPT zero-shot experiments results.
+| Unseen Answers | Unseen Questions | |||||||||||
+| Model | Acc. | F1 | BLEU | MET. | ROU. | BERT | Acc. | F1 | BLEU | MET. | ROU. | BERT |
+| Majority | 43.3 | 26.1 | 1.2 | 8.6 | 12.7 | 16.2 | 40.5 | 23.4 | 0.1 | 8.64 | 2.78 | 12.32 |
+| Fine-tuning experiments | ||||||||||||
+| 74.4 | 73.0 | 13.3 | 34.1 | 16.8 | 31.9 | 55.6 | 53.6 | 12.5 | 31.4 | 16.2 | 28.9 | |
+| 73.9 | 73.7 | 11.7 | 35.9 | 16.9 | 35.1 | 56.3 | 54.9 | 9.2 | 31.4 | 13.9 | 32.6 | |
+| 70.1 | 69.4 | 13.2 | 34.7 | 17.7 | 42.4 | 53.5 | 52.0 | 10.8 | 30.6 | 15.1 | 40.1 | |
+| 72.4 | 72.9 | 13.6 | 35.7 | 17.6 | 40.4 | 54.9 | 56.1 | 13.4 | 34.4 | 17.5 | 39.1 | |
+| 72.8 | 73.1 | 11.3 | 33.3 | 14.82 | 37.4 | 54.7 | 55.4 | 11.3 | 32.0 | 15.2 | 36.2 | |
+| 75.4 | 75.7 | 13.9 | 38.3 | 19.5 | 41.6 | 58.7 | 57.9 | 14.9 | 36.7 | 19.7 | 40.9 | |
+| Zero-shot experiments | ||||||||||||
+| ChatGPT-4o | 48.9 | 46.9 | 3.64 | 22.3 | 5.2 | 20.8 | - | - | - | - | - | - |
+| DeepSeek-R1 | 47.4 | 46.7 | 4.1 | 23.6 | 6.5 | 18.0 | - | - | - | - | - | - |
+Table 5 shows a majority baseline, fine-tuning experiments, and zero-shot experiments results.
 The majority baseline contains the most occurring label and feedback from the EngSAF train set.
@@
 You have answered the question correctly, covering all the required aspects.”.
-We conducted the ChatGPT experiment only on the unseen answers test set due to its zero-shot setting, as both unseen questions and answers are identical in the zero-shot setting.
-| Unseen Answers | Unseen Questions | |||||||||||
-| Model | Acc. | F1 | BLEU | MET. | ROU. | BERT | Acc. | F1 | BLEU | MET. | ROU. | BERT |
-| Majority | 43.3 | 26.1 | 1.2 | 8.6 | 12.7 | 16.2 | 40.5 | 23.4 | 0.1 | 8.64 | 2.78 | 12.32 |
-| 74.4 | 73.0 | 13.3 | 34.1 | 16.8 | 31.9 | 55.6 | 53.6 | 12.5 | 31.4 | 16.2 | 28.9 | |
-| 73.9 | 73.7 | 11.7 | 35.9 | 16.9 | 35.1 | 56.3 | 54.9 | 9.2 | 31.4 | 13.9 | 32.6 | |
-| 72.8 | 73.1 | 11.3 | 33.3 | 14.82 | 37.4 | 54.7 | 55.4 | 11.3 | 32.0 | 15.2 | 36.2 | |
-| 75.4 | 75.7 | 13.9 | 38.3 | 19.5 | 41.6 | 58.7 | 57.9 | 14.9 | 36.7 | 19.7 | 40.9 | |
-| ChatGPT | 41.1 | 24.0 | 7.4 | 31 | 11.3 | 28.8 | - | - | - | - | - | - |
 5.1 Analysis
 In this section, we delve into the insights drawn from the performance results of experiments on the EngSAF dataset for the ASAG task.
-The evaluation metrics used for labels are accuracy and macro-averaged F1.
-To evaluate the feedback, we measure the Rouge-2 Post (2018), SCAREBLEU 777https://pypi.org/project/sacrebleu/, METEOR Banerjee and Lavie (2005) and BERTSCORE Zhang et al.
-(2019) scores.
-Table 5 shows that Mistral-7b significantly outperforms the majority baseline in both label and feedback metrics.
-However, a notable disparity exists in improvement between unseen questions and unseen answers.
-We observe a decrease of approximately 23% in accuracy for unseen questions compared to unseen answers.
-This underscores the challenge even fine-tuned models face when tasked with generalizing to new questions.
-Additionally, there is approximately 4% decrease in Bertscore for unseen questions compared to unseen answers, suggesting the necessity for new evaluation metrics that comprehensively assess text on the context level instead of the lexical level.
-Including questions in the input has led to a notable improvement in Bertscore, with a 10% increase for unseen answers and a 12% increase for unseen questions.
-Similarly, for label accuracy, we observe a 3.5% increase for unseen answers and a 7.3% increase for unseen questions in the test set.
-Technical Appendix includes qualitative analysis of test examples.
-Overall, the experiment suggests including questions for improvement in the predicted feedback by the model.
-The ChatGPT zero-shot experiment yields the lowest accuracy at 41.1%, even falling below that of the majority baseline, highlighting the complexity of the ASAG task and struggle of large language models like ChatGPT struggle in a zero-shot setting.
-Another contributing factor to this performance disparity could be the inherent bias present in teacher grading, which models can only learn when trained on such data.
-Interestingly, the smaller Mistral 7B model outperformed the LLaMA-2-13B model for ASAG, demonstrating the potential of more compact models in this task.
-6 Deployment
+Following Filighera et at.’s [12] evaluation methodology, we measure the accuracy and macro-averaged F1 for output labels and the ROUGE-2 [26], SACREBLEU888https://pypi.org/project/sacrebleu/, METEOR [6] and BERTScore [37] for the feedback part.
+For all metrics, a higher score indicates better results.
+As per Table 5, Mistral-7b significantly outperforms the majority baseline in both output label and feedback metrics.
+However, we observe a significant performance gap between unseen answers and unseen questions, with accuracy dropping by approximately 23% for unseen questions and by 4% in BERTScore, suggesting the necessity for new evaluation metrics that comprehensively assess text on the context level instead of the lexical level.
+The above two observations highlight the challenge that even fine-tuned models face when generalizing to new questions or domains (UQ).
+In the case of unseen answers (UA), other responses to the same question are present in the training data, making generalization easier compared to entirely new questions.
+A similar pattern can be observed for other fine-tuned models.
+Notably, including questions in the input significantly improves accuracy, F1 score, and BERTScore.
+For example, in our experiments with Llama-2 and Mistral-7B, incorporating the question led to a 10% improvement in BERTScore for unseen answers and a 12% improvement for unseen questions.
+Our zero-shot experiments with ChatGPT and DeepSeek-R1 yield an accuracy of 48.9% and 47.4%, respectively, falling below the fine-tuning baselines, highlighting the complexity of the ASAG task.
+We observe a similar performance gap in the feedback evaluation part, suggesting that fine-tuning plays a crucial role in capturing complex grading patterns.
+Another contributing factor to this performance disparity of zero-shot experiments in comparison to fine-tuned experiments could be the inherent bias present in teacher grading, which models can only learn when trained on such data.
+6 Real-World Deployment
+For real-world deployment, we integrated our best-performing fine-tuned ASAG model for automatic evaluation of an end-semester exam of a course at a reputed institute.
+The end-semester examination included two short-answer questions, each accompanied by the instructor’s correct/reference answer.
+We randomly sampled 25 student answers for each question and used our fine-tuned model to predict the output label and feedback/ explanation for the predicted output.
+To ascertain the quality of the generated output, we employed three Subject Matter Experts (SMEs) to evaluate the predicted output and feedback.
+The SMEs were PhD students who were teaching assistants (TAs) of that course.
+Each predicted output label was evaluated based on its correctness, where the subject matter expert checks whether the predicted output label is correct or not.
+Each predicted feedback is analyzed in terms of Feedback Quality/Correctness (FC) and Emotional Impact (EI) as discussed in the Quality Estimate section 3.4.
+Each aspect is scored on a scale (1-5), with a high score indicating a better response.
 | Aspect | Avg. Score |
 | Output Label Accuracy | 92.5% |
-| Feedback Quality/Correctness | 4.5 |
-| Emotional Impact | 4.9 |
-For the real-world deployment, our fine-tuned ASAG model was integrated into an end-semester exam on ET 623 (Learning Analytics Course) at IIT Bombay for the 2024 academic year.
-The setup included the students enrolled in the course who consented to participate in the experiment.
-The end-sem exam includes 2 short-answer questions, each accompanied by the instructor’s correct/reference answer.
-We randomly sampled 25 student answers for each question and used our fine-tuned ASAG with Feedback model to predict the output label and feedback/ explanation for the predicted output.
-Each predicted output label was evaluated based on its correctness, where the subject matter expert checks whether the predicted output label is correct or not.
-Each predicted feedback is analyzed in terms of Feedback Correctness/ Accuracy and Emotional Impact as discussed in the Quality Estimate section 3.5.
-Each aspect is scored on a scale (1-5), with a high score indicating a better response.
-Table 6 presents the average scores across all the designed aspects used to measure the reliability of the proposed dataset.
-Upon examining the table, it is evident that the subject matter evaluation scores greater than 4.5 for both Feedback Quality/Correctness and the Emotional Impact aspect.
+| Feedback Quality/Correctness (FC) | 4.5 |
+| Emotional Impact (EI) | 4.9 |
+Table 6 presents the average scores across all the aspects used to measure the reliability of our ASAG model.
+Upon examining the table, it is evident that the subject matter evaluation scores are greater than 4.5 for both FC and the EI aspect.
 This demonstrates the reliability and effectiveness of the ASAG model in real-world scenarios.
 Additionally, the model’s predicted output label achieved an accuracy of 92.5%, further showcasing its performance and reliability.
-Technical Appendix contains more details about deployment.
-7 Summary, Conclusion, and Future Work
-This study presents a novel and extensive dataset for grading short answers across multiple engineering domains.
-Our dataset addresses the need for a standardized evaluation platform by covering diverse subjects and a wide range of short answers.
-We meticulously curated and annotated this dataset to ensure its quality and applicability.
-Using the Label Aware Synthetic Feedback Generation (LASFG) strategy we have synthetically incorporated feedback in the EngSAF dataset.
-Through experimentation, we benchmarked the dataset’s performance using different LLM’s fine-tuning and chatGPT zero shot experiments.
-In conclusion, our research endeavors culminated in successfully creating and evaluating a cutting-edge multi-domain short answer grading dataset.
-The dataset’s diverse content and meticulous annotations provide a solid platform for training and assessing grading models across various subject domains.
-Our benchmarking experiments showcased promising results, indicating that our dataset has the potential to significantly improve the quality and reliability of ASAG systems.
-Additionally, we deployed the ASAG system in a real-world setting for an IITB course exam, demonstrating its practical applicability.
-This work marks a crucial step forward in educational technology, equipping educators and researchers with a valuable resource to boost advancements in automated assessment techniques.
-While this research presents a novel dataset for effective short answer grading along with feedback, there are several possibilities for future investigation.
-One area of focus could be the refinement and expansion of our dataset to include more nuanced and complex short answers, enabling us to accommodate a broader range of grading scenarios.
-Baseline figures could be improved by expansion or augmentation of the dataset.
-External knowledge inclusion through knowledge graphs could be explored in this direction.
-8 Limitations
-Our study, titled “I understand why I got this grade": Automatic Short Answer Grading with Feedback" has provided valuable insights into the multi-domain ASAG.
-However, it’s important to acknowledge certain limitations.
-The synthetic nature of the generated feedback may introduce biases or might contain inconsistencies, factual inaccuracies, or overly generic content, that could impact the implications of the results.
-One key finding from table 3 is an imbalance in output labels.
-Specifically, Output label Zero (0) is less frequent compared to labels One (1) and Two (2).
-This imbalance could impact the performance of automatic grading models trained on this dataset.
-Despite our efforts to curate diverse short-answer responses across multiple engineering domains, the dataset’s overall size might limit the complexity and depth of models that can be trained on it.
-Variations in data distribution, language patterns, and task requirements across different domains and contexts may impact the application of our results to real-world scenarios.
-While our dataset’s strength lies in its coverage of multiple engineering domains, it’s important to note that the range of domains might not be exhaustive.
-Researchers focusing on specific domains or aiming to enhance domain-specific performance may need to consider domain-specific or subject-specific fine-tuning or gather additional data.
-9 Ethical Statement
-The EngSAF dataset contains questions, reference answers, and student answers along with the output label assigned for each student answer.
-This dataset is taken from actual exams and quizzes conducted at the university.
-We are committed to protecting those student’s privacy and anonymity.
-Hence, no student-specific information is disclosed within the dataset.
-The process of collecting and curating this dataset adhered to the ethical norms of informed consent.
-Prior to any contribution, instructors of each course were provided details concerning the data collection’s intent, the utilization of their questions and responses, and any potential implications.
-While the ASAG models created with this dataset offer great potential in educational assessment, they must be used properly and ethically.
-As with any AI technology, its impact on education must be regularly monitored and evaluated.
-Ensuring an appropriate balance between the advantages derived from automation and safeguarding the educational experience and human judgement is of utmost importance.
+By providing accurate labels and feedback at scale, our approach helps reduce the manual workload for teachers and instructors, making the grading process more efficient and scalable.
+7 Conclusion
+This work introduces a novel and extensive dataset for automatic short answer grading (ASAG) across multiple engineering domains.
+Addressing the need for a standardized evaluation platform, our dataset encompasses diverse subjects and a wide range of short-answer responses.
+We meticulously curated and annotated the dataset to ensure its quality, consistency, and applicability to real-world grading scenarios.
+We propose Label-Aware Synthetic Feedback Generation (LASFG)—a method that enables the transformation of any traditional ASAG dataset into one that incorporates meaningful synthetic feedback.
+Additionally, we benchmark our dataset using fine-tuned large language models (LLMs) and conduct zero-shot evaluations, providing a comprehensive assessment of its effectiveness.
+Furthermore, we deployed our ASAG system in a real-world course exam at a prestigious institution, showcasing its practical utility in educational settings.
+Human evaluation showcases the effectiveness of the ASAG system by achieving scores more than 4.5 out of 5 on Feedback Quality/Correctness (FC) and Emotional Impact (EI).
+Additionally, the model demonstrates its ability to correctly detect an answer with an accuracy of 92.5%.
+This work represents a significant milestone in educational technology, offering educators and researchers a valuable tool to drive innovation in automated assessment methods.
+Future extensions of this work could focus on expanding the dataset to include more complex short-answer responses, enabling broader grading scenarios.
+Additionally, incorporating the evaluation of diagrams and images within student responses could further enhance the system’s capabilities, paving the way for more comprehensive automated assessment solutions.